(2 points) Determine the scalar product of A= 6.0i + 4.0j - 2.0k andB = 5.0i - 6.0 j- 3.0 k.
Answer #1

The scalar product or dot product of any two vectors is the product of magnitudes of the two vectors and the cosine of the angle between them.

If \underset{A}{\rightarrow} and \underset{B}{\rightarrow} are the two vectors include with an angle \theta,then the scalar product of the two vectors is ,

\underset{A}{\rightarrow}.\underset{B}{\rightarrow}= AB COS\theta

Scalar product can be written in terms of magnitude of their components as,

\underset{A}{\rightarrow} = 6 i + 4 j - 2k

\underset{B}{\rightarrow} = 5i -6j -3k

\underset{A}{\rightarrow}.\underset{B}{\rightarrow} = ( 6* 5)( i.i) +( 4* -6)(j.j) +(-2 * -3)(k.k)

= 30 - 24 + 6

= 12 units

since i.i = j.j = k.k =1